OpenStudy (amberlykhan):

Find the largest factor of 1260 which is not divisible by 6. Explain.

OpenStudy (anonymous):

do you have a prime factorization for 1260?

OpenStudy (anonymous):

Numbers divisible by 6 are divisible by both 2 and 3, so once you have the prime factors, multiply all of them except the 2s to get the largest number not divisible by 6.

OpenStudy (anonymous):

\[1260=2^2 *3^2* 5 *7\]

OpenStudy (anonymous):

315

OpenStudy (anonymous):

3^2∗5∗7

OpenStudy (amberlykhan):

Thanks! @peachpi and @robtobey. But why would I multiply the 3s if you stated that 6 is divisible by 2 and 3? Kinda confused about why you canceled out the 2s to get the answer.

OpenStudy (loser66):

Because 2 and 3 cannot be at the expression at the same time. If they exist at the same time, then the number will be divisible by 6

OpenStudy (amberlykhan):

So why cancel out 2 but not 3? If only one can be present, cant it be 3 also? @Loser66

OpenStudy (loser66):

Because when you cancel 3, the number is not divisible by 6 but it is NOT the largest one.

OpenStudy (loser66):

\(3^2*5*7 = 325 > 2^2*5*7 = 140\)

OpenStudy (amberlykhan):

Ohhhhhh..... Thanks! I FINALLY get it. :) @Loser66

OpenStudy (loser66):

np